OpenStudy (calculusxy):

If 30% of m is 40, what is 15% of m?

OpenStudy (gabebae):

I'll help you :')

OpenStudy (calculusxy):

Thanks!

OpenStudy (gabebae):

Look closely at it You're missing a huge part of it :P What is half of 30%

OpenStudy (calculusxy):

15%. I was actually thinking about this but in a proportional way \[\frac{ 30 }{ 40 } \rightarrow \frac{ 3 }{ 4 } = \frac{ 15 }{ x } \implies x = 20\]

OpenStudy (gabebae):

That's what I did as well

OpenStudy (freckles):

you could setup the equations replace is with = replace of with * <--that is a times symbol. \[.3 m=40 \text{ then what is } .15 m=?\]

OpenStudy (calculusxy):

Cool!

OpenStudy (gabebae):

I got 20 too

OpenStudy (calculusxy):

Thanks @GabeBae

OpenStudy (gabebae):

You're welcome